# Practical aspects for the creation of an audio dataset from field recordings with optimized labeling budget with AI-assisted strategy

### v2

@@ 4 Conclusion
This paper establishes a complete framework for the creation of an audio database while optimizing the labeling budget in high-volume unlabeled data environments. This framework could be extremely useful when crowdsourcing is to be avoided to keep insight during the labeling process. The use case presented has been carried out in a commercial-industrial environment such as the port of Valencia (Spain). At the time of drafting this paper, 6540 10-second audios have been tagged with an effort of 52 hours per labeler having a pool of 5 labelers divided into two groups of 2 and 3 members. A dataset is expected to be released soon as one of the outcomes of this work.
Broader Impact Statement
The work proposes a framework to be used for Machine Listening projects in scenarios with constrained tagging budgets. General aspects of data collection, the AL-based labeling with AI-assisted techniques and the storage of the information obtained from the labeling process are presented. In Machine Listening projects it is common that the largest and best-known sounding databases are either not suited to the particular problems of the project or cannot be used for commercial purposes. That is why a framework such as the one presented here should become standard practice in any AI-based machine listening project.
@@
The participation of Javier Naranjo-Alcazar, Jordi Grau-Haro and Pedro Zuccarello in this research was funded by the Valencian Institute for Business Competitiveness (IVACE) and the FEDER funds by means of project Soroll-IA2 (IMDEEA/2023/91).
Appendix A Ontology
The labels initially used for labeling correspond to a set of 527 tags derived from the Audioset dataset (Gemmeke et al. (2017)). These classes were used to train the PANNs network (Kong et al. (2020)). Due to the proposals of the taggers in the weekly meetings (process explained in Section 3) nine additional tags were incorporated to encompass the scenarios observed in the node recordings. These newly added classes are:
  • Car door closing
  • Seagull
  • Metal noises
  • Disc grinder
  • Construction noises
  • Blower
  • Packing tape, duct tape
  • Flange
  • Crane
The number of tags is anticipated to increase as additional audio recordings are labeled.
The current tag configuration permits high granularity in labeling. This configuration enables the execution of various training processes depending on the desired level of granularity. Consequently, similar tags or tags associated with the same parent node, as defined by the Audioset structure Gemmeke et al. (2017), can be grouped accordingly.

### v3

@@ 4 Conclusion
This paper establishes a complete framework for the creation of an audio database while optimizing the labeling budget in high-volume unlabeled data environments. This framework could be extremely useful when crowdsourcing is to be avoided to keep insight during the labeling process. The use case presented has been carried out in a commercial-industrial environment such as the port of Valencia (Spain). At the time of drafting this paper, 6540 10-second audios have been tagged with an effort of 52 hours per labeler having a pool of 5 labelers divided into two groups of 2 and 3 members.
As future work, continuing with the labeling process is expected to further increase the labeled database. Regarding the proposed Active Learning pipeline, the aim is to work on the algorithm itself, being able to propose new classifiers without label propagation because this technique could imply a problem once a large volume of labeled data is available. Finally, in the near future, it is intended to make public a subset of the labeled data for research use.
@@
The participation of Javier Naranjo-Alcazar, Jordi Grau-Haro and Pedro Zuccarello in this research was funded by the Valencian Institute for Business Competitiveness (IVACE) and the FEDER funds by means of project Soroll-IA2 (IMDEEA/2023/91). The research carried out for this publication has been partially funded by the project STARRING-NEURO (PID2022-137048OA-C44) funded by the Ministry of Science, Innovation and Universities of Spain and the European Union.